OpenStudy (anonymous):

solve the system 3y=-x+11 x+4y=14

OpenStudy (anonymous):

(2,3)

OpenStudy (anonymous):

x+3y=11 -x-4y=-14 0-y=-3 y=3

OpenStudy (anonymous):

x=2, y=3

OpenStudy (anonymous):

i am being annoying sorry. just trying to see if i can be fast. i will stop

OpenStudy (anonymous):

thank you